A rope of uniform mass M and length L hangs from the branch of a tree.Find the tension at a distance x from the ground.

Grade:11

2 Answers

Aman Bansal
11 years ago

Dear Rahul,

mass of thr rope is m
length of the rope is L .
Then mass per unit length is μ = m/L
then tension at distance y  is T = m''g
m'' = mass of the part having length y = μy
then tension T =  μy*g
then speed of transverse wave is v = √T/μ =√μy*g /μ  = √(gy)
b) distance travelled by the transverse wave is s = L
   velocity of wave at the bottomis u = 0
   velocity of the wave at the top is v =√gL
  Then distance s = average velocity *time
      L =[ (v+u)/2]*t
        = [ (√gL)/2]*t
then time t = 2√g/L

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

Rahul Nath.R
11 years ago

But the answer to this question in the book is tension =0 at the height x from the ground.